OpenStudy (anonymous):

Find the equation in slope-intercept for of the line that has 2 and passes through A(-8,7)

OpenStudy (haseeb96):

y-y1=m(x-x1) Here (x1,y1)=(-8,7) m=2 Now just put them into the above formula and get the answer

OpenStudy (haseeb96):

y-7=2(x+8) y-7=2x+16 y=2x+8+7 y=2x+15 This is your answer in slope intercept form

OpenStudy (anonymous):

Thank you! But the math is wrong I think because I think it's supposed to be y=2x+16+7 which would give you y=2x+23

OpenStudy (haseeb96):

Yes sorry that was my mistake.